A 0.271-g sample of a gaseous hydrocarbon (CxHy) occupies 226 mL at 22.0 ºC at 735 torr. Calculate the molar mass of this hydrocarbon in g/mol. Express your answer as a number (not scientific notation) without units; for full credit, the last digit of your answer should be within ±1 of the correct value, with an appropriate number of significant figures.

Molar mass: Answer          g/mol.

Complete combustion of 1.532 g of this gaseous hydrocarbon (CxHy) yields 4.484 g of carbon dioxide and 2.754 g of water. Calculate the moles of C and H in this amount of the hydrocarbon. Express your answers as a number (not scientific notation) without units; for full credit, the last digit of your answer should be within ±2 of the correct value, with an appropriate number of significant figures.

Answer moles C in CxHy

Answer moles H in CxHy

Determine the empirical and molecular formulas of this hydrocarbon. If a subscript in the formula is "1", enter it explicitly.

Empirical formula: CAnswerHAnswer.

Molecular formula: CAnswerHAnswer.
Answer #1

i 1 me = 103 226 mL = 2268103 L ó, Oe = 273 k. 22.0°e = 273+ 2205 295.0 .. I torr = 0.00 13 atm 735 tort = 0.967 atm using I:. mote = mass molar mass molar mass = mass = 0,271 g = 30g/mol 9,02 x 103g 1mol . Cuty + puty > xCO2 + y the moles of co2 =